OpenStudy (cicilybailey14):

I MEDAL!! The recommended dosage D in milligrams of a certain medicine depends on a person's body mass "w" in kilograms. The function rule D = 0.1w^2 + 5w describes the relationship of the dosage to body mass. Evaluate the function for a person who has a mass of 60 kilograms.

OpenStudy (anonymous):

0.1 *60^2 + 5*60=

OpenStudy (anonymous):

0.1*3600 +300 = 360+300 = 660

OpenStudy (cicilybailey14):

thx!!